The width of a rectangle is three less than the lenght. if the perimeter of the rectangle is 30 yards finr the width.
Answer by checkley77(12844)

W=L+3
P=2L+2W
30=2L+2(L+3)
30=2L+2L+6
30=4L+6
4L=30-6
4L=24
L=24/4
L=6 ANS. FOR THE LENGTH.
W=6+3=9 ANS. FOR THE WIDTH.
PROOF:
30=2*9+2*6
30=18+12
30=30